Your weekly paycheck is 10 percent less than your coworker's. Your two paychecks total 640. Find the amount of each paycheck.
Answer by bucky(2189)

Presume that the unknown amount that your coworker gets in the weekly paycheck is represented
by W. Since you make 10 percent less than that coworker, you make 90 percent (a factor of
0.9 times) of what that coworker does. So the combined amount that the two of you make can be
represented as:
.
W + 0.9W
.
and the problem tells you that this sum is 640. So you can write the equation:
.
W + 0.9W = 640
.
Add the two terms on the left side and you get:
.
1.9W = 640
.
Solve for W by dividing both sides of this equation by 1.9 and you have:
.
W = 640/1.9 = 336.84 (rounded off to the nearest hundredth)
.
That is what your coworker makes. Since you make 90 percent of that, the amount you make is:
.
0.9W = 0.9(336.84) = 303.16 (rounded off to the nearest hundredth)
.
Don't forget to label you answer as euros, yen, rials, pounds, dollars, pesos or whatever
currency you are working with.
.
You can check this by adding the two checks together (336.84 + 303.16) and you will see that
they do total to 640.
.
Hope this helps you to understand the problem.
.